complete the square, and find the roots of the quadratic equation.
x2 + 6x -16= 0

Answer by nerdybill(7384)

x^2 + 6x -16= 0
x^2 + 6x = 16
take half 'b', square it and add to both sides:
x^2 + 6x + 9 = 16 + 9
(x+3)(x+3) = 25
(x+3)^2 = 25
x+3 = +-5
x = -3+-5
x = {-8, 2}